r the work you have completed so far. It does not indicat Norwall Companys budgeted variable manufacturing overhead cost is $1.05 per machine-hour and its budgeted fixed manufacturing overhead is $31,382 per month. a. The denominator activity of 8,840 machine-hours is used to compute the predetermined overheqd rate. so b. At a denominator activity of 8,840 machine-hours, the company should produce 3,400 units of product. c. The companys actual operating results were: Number of units produced Actual machine-hours Actual variable manufacturing overhead cost Actual fixed manufacturing overhead cost 4,360 s 11,964 s 30,000 the predetermined overhead rate and break it down into variable and fixed cost elements. (Round your answers to 2 decimal places.) 3. Compute the variable overhead rate and e Input all amounts as positive values. Round your intermediate calculations and final answers to 2 decimal places.) the variable overhead rate and efficiency variances and the fixed overhead budget and volume variances.(ndicate able, and None for no effect (i.e., zero variance). F for 460per MH 1.05per MH 355per MH < Prev 50, 7111 Next > I just need number 3
2. Compute the standard hours allowed for the actual production 3. Compute the variable overhead rate and efficiency variances and the fixed o the effect of each variance by selecting F for favorable, U for unfavorable, and N Input all amounts as positive values. Round your intermediate calculations and final ans verhead buc Answer is not complete. Predetermined overhead rate Variable element Fixed element 4.60。1 per MH 1.05per MH 3.55per MH 11,336MHs 1. 2. Standard hours allowed for the actual production 3. Variable overhead rate variance Variable overhead efficiency variance Fixed overhead budget variance Fixed overhead volume variance Prev
Answer #1

Answer:

Predetermined overhead rate

4.6

Variable element

1.05

Fixed element (4.6-1.05)

3.55

Standard Hours allowed for actual Production

11336

Variable Overhead rate variance

1495.5

U

Variable Overhead efficiency variance

1434.3

F

Fixed Overhead Budget Variance

1382

F

Fixed Overhead Volume Variance

8860.8

F

Working notes for the above answer is as under

1

Predetermined overhead rate

={(1.05*8840)+31382} /8840

=40664/8840

=4.6

2

Standard Hours allowed for actual Production

=Actual Units product x standard Hour Required for One unit

=4360 x (8840/3400)

=4360 x 2.6

=11,336 hour

3

Variable Overhead rate variance

=(AH x AR)- (AH x SR)

=(11964)- (9970*1.05)

=11964-10468.5

=1495.5 U

4

Variable Overhead efficiency variance

=Standard rate x (actual hour- standard Hour)

=1.05 x (9970-11336)

=1434.3 Favorable

5

Fixed Overhead Budget Variance

=Actual Fixed overhead cost-Budgeted Fixed overhead cost

=30,000-31382

=1382 favorable

6

Fixed Overhead Volume Variance

=Fixed Portion of Predetermined Overhead x (Denominator Hour - Standered Hour)

=3.55 x (8840-11336)

=3.55x 2496

=8860.8 Favorable